OpenStudy (anonymous):

Percent of increase 2 hoodies to 32 hoodies

OpenStudy (anonymous):

Thirty two is sixteen time two, so the increase is fifteen times, or 1500%

OpenStudy (anonymous):

32-2= -30 absolute value = 30. Now 30/2 =15 *100 = 1,500%